OpenStudy (nickymail48):

how to solve -2y-12=6y+36

OpenStudy (anonymous):

First try to isolate like terms on left side & all other on right hand side that is move 6y to left hand side by subtracting 6y from both sides & move -12 to right hand side by adding 12 to both sides, So what do you get ?

OpenStudy (solomonzelman):

Yes, to factor both sides completely.

OpenStudy (nickymail48):

-8y+24?

OpenStudy (anonymous):

Please check it at left side we get -8y 7 right hand side 36+12

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf-2y-12=6y+36 }\) \(\LARGE\color{blue}{ \bf-2(y+6)=6(y+6) }\) so what do you think now ?

OpenStudy (solomonzelman):

@tgawade, you didn't ask the question, so let the asker do it. Also you are not correct. Now, @nickymail48 , based on what I posted can you come up with the correct answer?

OpenStudy (nickymail48):

3

OpenStudy (anonymous):

We have -2y - 12 = 6y + 36 So first we isolate like terms to left hand side to move 6y to left hand side we subtract 6y from both sides so we get -2y -6y -12 = 6y - 6y +36 therefore -8y -12 = 36 now to move -12 to right hand side we add12 on both sides -8y -12 + 12 = 36 + 12 -8y = 48 now to get rid of -8 from left side we divide both sides by -8 so we will get -8y/-8 = 48/-8 y = -6

OpenStudy (anonymous):

have you understood ??

OpenStudy (nickymail48):

yes

OpenStudy (nickymail48):

thank you

OpenStudy (anonymous):

you are welcome :)